Find all real number solutions to the following equations. Please answer in radians.

a. sin(−?) = − 1/2

b. tan 2? = −√3
Answer #1

NEI la sin (-2) = -1/ a sinn = 1 & cina - 4 = sinto We know it sino a sina then a = MT + (-11 na so x = MT + 1-1M I ar na I c